\begin{thebibliography}{28}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Artetxe and Schwenk(2019)}]{Artetxe2019MassivelyMS}
Mikel Artetxe and Holger Schwenk. 2019.
\newblock Massively multilingual sentence embeddings for zero-shot cross-lingual transfer and beyond.
\newblock \emph{Transactions of the Association for Computational Linguistics}, 7:597--610.

\bibitem[{Campos et~al.(2016)Campos, Nguyen, Rosenberg, Song, Gao, Tiwary, Majumder, Deng, and Mitra}]{Campos2016MSMA}
Daniel~Fernando Campos, Tri Nguyen, Mir Rosenberg, Xia Song, Jianfeng Gao, Saurabh Tiwary, Rangan Majumder, Li~Deng, and Bhaskar Mitra. 2016.
\newblock \href {https://arxiv.org/abs/1611.09268} {Ms marco: A human generated machine reading comprehension dataset}.
\newblock \emph{ArXiv preprint}, abs/1611.09268.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek, Guzm{\'a}n, Grave, Ott, Zettlemoyer, and Stoyanov}]{conneau2020unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzm{\'a}n, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.747} {Unsupervised cross-lingual representation learning at scale}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 8440--8451, Online. Association for Computational Linguistics.

\bibitem[{Costa-juss{\`a} et~al.(2022)Costa-juss{\`a}, Cross, {\c{C}}elebi, Elbayad, Heafield, Heffernan, Kalbassi, Lam, Licht, Maillard et~al.}]{costa2022no}
Marta~R Costa-juss{\`a}, James Cross, Onur {\c{C}}elebi, Maha Elbayad, Kenneth Heafield, Kevin Heffernan, Elahe Kalbassi, Janice Lam, Daniel Licht, Jean Maillard, et~al. 2022.
\newblock No language left behind: Scaling human-centered machine translation.
\newblock \emph{arXiv preprint arXiv:2207.04672}.

\bibitem[{DataCanary et~al.(2017)DataCanary, hilfialkaff, Jiang, Risdal, Dandekar, and tomtung}]{quora-question-pairs}
DataCanary, hilfialkaff, Lili Jiang, Meg Risdal, Nikhil Dandekar, and tomtung. 2017.
\newblock \href {https://kaggle.com/competitions/quora-question-pairs} {Quora question pairs}.

\bibitem[{Fan et~al.(2019)Fan, Jernite, Perez, Grangier, Weston, and Auli}]{fan2019eli5}
Angela Fan, Yacine Jernite, Ethan Perez, David Grangier, Jason Weston, and Michael Auli. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1346} {{ELI}5: Long form question answering}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics}, pages 3558--3567, Florence, Italy. Association for Computational Linguistics.

\bibitem[{Feng et~al.(2022)Feng, Yang, Cer, Arivazhagan, and Wang}]{feng2022language}
Fangxiaoyu Feng, Yinfei Yang, Daniel Cer, Naveen Arivazhagan, and Wei Wang. 2022.
\newblock Language-agnostic bert sentence embedding.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 878--891.

\bibitem[{Gao et~al.(2021)Gao, Yao, and Chen}]{Gao2021SimCSESC}
Tianyu Gao, Xingcheng Yao, and Danqi Chen. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.emnlp-main.552} {{S}im{CSE}: Simple contrastive learning of sentence embeddings}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing}, pages 6894--6910, Online and Punta Cana, Dominican Republic. Association for Computational Linguistics.

\bibitem[{Izacard et~al.(2021)Izacard, Caron, Hosseini, Riedel, Bojanowski, Joulin, and Grave}]{Izacard2021TowardsUD}
Gautier Izacard, Mathilde Caron, Lucas Hosseini, Sebastian Riedel, Piotr Bojanowski, Armand Joulin, and Edouard Grave. 2021.
\newblock \href {https://arxiv.org/abs/2112.09118} {Towards unsupervised dense information retrieval with contrastive learning}.
\newblock \emph{ArXiv preprint}, abs/2112.09118.

\bibitem[{Karpukhin et~al.(2020)Karpukhin, Oguz, Min, Lewis, Wu, Edunov, Chen, and Yih}]{Karpukhin2020DensePR}
Vladimir Karpukhin, Barlas Oguz, Sewon Min, Patrick Lewis, Ledell Wu, Sergey Edunov, Danqi Chen, and Wen-tau Yih. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.550} {Dense passage retrieval for open-domain question answering}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP)}, pages 6769--6781, Online. Association for Computational Linguistics.

\bibitem[{Lo et~al.(2020)Lo, Wang, Neumann, Kinney, and Weld}]{lo2020s2orc}
Kyle Lo, Lucy~Lu Wang, Mark Neumann, Rodney Kinney, and Daniel~S Weld. 2020.
\newblock S2orc: The semantic scholar open research corpus.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 4969--4983.

\bibitem[{Muennighoff et~al.(2023)Muennighoff, Tazi, Magne, and Reimers}]{muennighoff2023mteb}
Niklas Muennighoff, Nouamane Tazi, Loic Magne, and Nils Reimers. 2023.
\newblock \href {https://aclanthology.org/2023.eacl-main.148} {{MTEB}: Massive text embedding benchmark}.
\newblock In \emph{Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics}, pages 2014--2037, Dubrovnik, Croatia. Association for Computational Linguistics.

\bibitem[{Muennighoff et~al.(2022)Muennighoff, Wang, Sutawika, Roberts, Biderman, Scao, Bari, Shen, Yong, Schoelkopf et~al.}]{muennighoff2022crosslingual}
Niklas Muennighoff, Thomas Wang, Lintang Sutawika, Adam Roberts, Stella Biderman, Teven~Le Scao, M~Saiful Bari, Sheng Shen, Zheng-Xin Yong, Hailey Schoelkopf, et~al. 2022.
\newblock Crosslingual generalization through multitask finetuning.
\newblock \emph{arXiv preprint arXiv:2211.01786}.

\bibitem[{Ni et~al.(2022{\natexlab{a}})Ni, Hernandez~Abrego, Constant, Ma, Hall, Cer, and Yang}]{ni2022sentence}
Jianmo Ni, Gustavo Hernandez~Abrego, Noah Constant, Ji~Ma, Keith Hall, Daniel Cer, and Yinfei Yang. 2022{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2022.findings-acl.146} {Sentence-t5: Scalable sentence encoders from pre-trained text-to-text models}.
\newblock In \emph{Findings of the Association for Computational Linguistics: ACL 2022}, pages 1864--1874, Dublin, Ireland. Association for Computational Linguistics.

\bibitem[{Ni et~al.(2022{\natexlab{b}})Ni, Qu, Lu, Dai, Hernandez~Abrego, Ma, Zhao, Luan, Hall, Chang, and Yang}]{Ni2021LargeDE}
Jianmo Ni, Chen Qu, Jing Lu, Zhuyun Dai, Gustavo Hernandez~Abrego, Ji~Ma, Vincent Zhao, Yi~Luan, Keith Hall, Ming-Wei Chang, and Yinfei Yang. 2022{\natexlab{b}}.
\newblock \href {https://aclanthology.org/2022.emnlp-main.669} {Large dual encoders are generalizable retrievers}.
\newblock In \emph{Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing}, pages 9844--9855, Abu Dhabi, United Arab Emirates. Association for Computational Linguistics.

\bibitem[{OpenAI(2023)}]{OpenAI2023GPT4TR}
OpenAI. 2023.
\newblock \href {https://arxiv.org/abs/2303.08774} {Gpt-4 technical report}.
\newblock \emph{ArXiv preprint}, abs/2303.08774.

\bibitem[{Qiu et~al.(2022)Qiu, Li, Qu, Chen, She, Liu, Wu, and Wang}]{qiu2022dureader}
Yifu Qiu, Hongyu Li, Yingqi Qu, Ying Chen, QiaoQiao She, Jing Liu, Hua Wu, and Haifeng Wang. 2022.
\newblock \href {https://aclanthology.org/2022.emnlp-main.357} {{D}u{R}eader-retrieval: A large-scale {C}hinese benchmark for passage retrieval from web search engine}.
\newblock In \emph{Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing}, pages 5326--5338, Abu Dhabi, United Arab Emirates. Association for Computational Linguistics.

\bibitem[{Reimers and Gurevych(2019)}]{Reimers2019SentenceBERTSE}
Nils Reimers and Iryna Gurevych. 2019.
\newblock \href {https://doi.org/10.18653/v1/D19-1410} {Sentence-{BERT}: Sentence embeddings using {S}iamese {BERT}-networks}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP)}, pages 3982--3992, Hong Kong, China. Association for Computational Linguistics.

\bibitem[{Thorne et~al.(2018)Thorne, Vlachos, Christodoulopoulos, and Mittal}]{Thorne2018FEVERAL}
James Thorne, Andreas Vlachos, Christos Christodoulopoulos, and Arpit Mittal. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-1074} {{FEVER}: a large-scale dataset for fact extraction and {VER}ification}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers)}, pages 809--819, New Orleans, Louisiana. Association for Computational Linguistics.

\bibitem[{Wang et~al.(2022)Wang, Yang, Huang, Jiao, Yang, Jiang, Majumder, and Wei}]{wang2022text}
Liang Wang, Nan Yang, Xiaolong Huang, Binxing Jiao, Linjun Yang, Daxin Jiang, Rangan Majumder, and Furu Wei. 2022.
\newblock \href {https://arxiv.org/abs/2212.03533} {Text embeddings by weakly-supervised contrastive pre-training}.
\newblock \emph{ArXiv preprint}, abs/2212.03533.

\bibitem[{Wang et~al.(2023)Wang, Yang, Huang, Yang, Majumder, and Wei}]{wang2023improving}
Liang Wang, Nan Yang, Xiaolong Huang, Linjun Yang, Rangan Majumder, and Furu Wei. 2023.
\newblock Improving text embeddings with large language models.
\newblock \emph{arXiv preprint arXiv:2401.00368}.

\bibitem[{Wang et~al.(2021)Wang, Bao, Huang, Dong, and Wei}]{wang2021minilmv2}
Wenhui Wang, Hangbo Bao, Shaohan Huang, Li~Dong, and Furu Wei. 2021.
\newblock Minilmv2: Multi-head self-attention relation distillation for compressing pretrained transformers.
\newblock In \emph{Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021}, pages 2140--2151.

\bibitem[{Xiao et~al.(2023)Xiao, Liu, Zhang, and Muennighof}]{xiao2023c}
Shitao Xiao, Zheng Liu, Peitian Zhang, and Niklas Muennighof. 2023.
\newblock \href {https://arxiv.org/abs/2309.07597} {C-pack: Packaged resources to advance general chinese embedding}.
\newblock \emph{ArXiv preprint}, abs/2309.07597.

\bibitem[{Xue et~al.(2021)Xue, Constant, Roberts, Kale, Al-Rfou, Siddhant, Barua, and Raffel}]{xue2021mt5}
Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2021.
\newblock mt5: A massively multilingual pre-trained text-to-text transformer.
\newblock In \emph{Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies}, pages 483--498.

\bibitem[{Yang et~al.(2018)Yang, Qi, Zhang, Bengio, Cohen, Salakhutdinov, and Manning}]{yang2018hotpotqa}
Zhilin Yang, Peng Qi, Saizheng Zhang, Yoshua Bengio, William Cohen, Ruslan Salakhutdinov, and Christopher~D. Manning. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1259} {{H}otpot{QA}: A dataset for diverse, explainable multi-hop question answering}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing}, pages 2369--2380, Brussels, Belgium. Association for Computational Linguistics.

\bibitem[{Zhang et~al.(2021)Zhang, Ma, Shi, and Lin}]{zhang2021mr}
Xinyu Zhang, Xueguang Ma, Peng Shi, and Jimmy Lin. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.mrl-1.12} {Mr. {T}y{D}i: A multi-lingual benchmark for dense retrieval}.
\newblock In \emph{Proceedings of the 1st Workshop on Multilingual Representation Learning}, pages 127--137, Punta Cana, Dominican Republic. Association for Computational Linguistics.

\bibitem[{Zhang et~al.(2023)Zhang, Thakur, Ogundepo, Kamalloo, Alfonso-Hermelo, Li, Liu, Rezagholizadeh, and Lin}]{Zhang2023MIRACLAM}
Xinyu~Crystina Zhang, Nandan Thakur, Odunayo Ogundepo, Ehsan Kamalloo, David Alfonso-Hermelo, Xiaoguang Li, Qun Liu, Mehdi Rezagholizadeh, and Jimmy Lin. 2023.
\newblock Miracl: A multilingual retrieval dataset covering 18 diverse languages.
\newblock \emph{Transactions of the Association for Computational Linguistics}, 11:1114--1131.

\bibitem[{Zweigenbaum et~al.(2018)Zweigenbaum, Sharoff, and Rapp}]{zweigenbaum2018overview}
Pierre Zweigenbaum, Serge Sharoff, and Reinhard Rapp. 2018.
\newblock Overview of the third bucc shared task: Spotting parallel sentences in comparable corpora.
\newblock In \emph{Proceedings of 11th Workshop on Building and Using Comparable Corpora}, pages 39--42.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[]{ACL2023}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}
\usepackage[hang,flushmargin]{footmisc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{booktabs}       
\usepackage{url}            
\usepackage{amsfonts}       
\usepackage{amsmath}
\usepackage{graphicx}
\usepackage{multirow}
\usepackage{pifont}
\usepackage{natbib}

\title{Multilingual E5 Text Embeddings: A Technical Report}

\author{Liang Wang,~Nan Yang,~Xiaolong Huang,
~Linjun Yang,~Rangan Majumder,~Furu Wei\\
Microsoft Corporation \\
\{wangliang,nanya,fuwei\}@microsoft.com
}

\begin{document}
\maketitle
\begin{abstract}
This technical report presents the training methodology and evaluation results of the open-source multilingual E5 text embedding models,
released in mid-2023.
Three embedding models of different sizes (small / base / large) are provided,
offering a balance between the inference efficiency and embedding quality.
The training procedure adheres to the English E5 model recipe,
involving contrastive pre-training on 1 billion multilingual text pairs,
followed by fine-tuning on a combination of labeled datasets.
Additionally,
we introduce a new instruction-tuned embedding model,
whose performance is on par with state-of-the-art, English-only models of similar sizes.
Information regarding the model release can be found at \url{https://github.com/microsoft/unilm/tree/master/e5}.
\end{abstract}

\section{Introduction}
Text embeddings serve as fundamental components in information retrieval systems
and retrieval-augmented language models.
Despite their significance,
most existing embedding models are trained exclusively on English text~\citep{Reimers2019SentenceBERTSE,Ni2021LargeDE,ni2022sentence},
thereby limiting their applicability in multilingual contexts.

In this technical report,
we present the multilingual E5 text embedding models (\emph{mE5-\{small / base / large\}}),
which extend the English E5 models~\citep{wang2022text}.
The training procedure adheres to the original two-stage methodology:
weakly-supervised contrastive pre-training on billions of text pairs,
followed by supervised fine-tuning on small quantity of high-quality labeled data.
We also release an instruction-tuned embedding model~\footnote{Here instructions refer to the natural language descriptions of the embedding tasks.}
\emph{mE5-large-instruct} by utilizing the synthetic data from ~\citet{wang2023improving}.
Instructions can better inform embedding models about the task at hand,
thereby enhancing the quality of the embeddings.

For model evaluation,
we first demonstrate that our multilingual embeddings exhibit competitive performance
on the English portion of the MTEB benchmark~\citep{muennighoff2023mteb},
and the instruction-tuned variant even surpasses strong English-only models of comparable sizes.
To showcase the multilingual capability of our models,
we also assess their performance on the MIRACL multilingual retrieval benchmark~\citep{Zhang2023MIRACLAM}
across $16$ languages and on Bitext mining~\citep{zweigenbaum2018overview,Artetxe2019MassivelyMS}
in over $100$ languages.

\section{Training Methodology}

\begin{table}[ht]
\centering
\scalebox{0.95}{\begin{tabular}{lc}
\hline
          & \# Sampled \\ \hline
Wikipedia &  150M    \\
mC4  &  160M  \\
Multilingual CC News   &  160M \\
NLLB  &   160M  \\
Reddit  &  160M   \\
S2ORC  &   50M   \\
Stackexchange   &  50M \\
xP3  &  80M   \\
Misc. SBERT Data   &   10M     \\ \hline
Total          &    $\sim$1B        \\ \hline
\end{tabular}}
\caption{Data mixture for contrastive pre-training.}
\label{tab:pretrain_data}
\end{table}

\noindent
\textbf{Weakly-supervised Contrastive Pre-training }
In the first stage,
we continually pre-train our model on a diverse mixture of multilingual text pairs
obtained from various sources as listed in Table~\ref{tab:pretrain_data}.
The models are trained with a large batch size $32k$ for a total of $30k$ steps,
which approximately goes over $\sim1$ billion text pairs.
We employ the standard InfoNCE contrastive loss with only in-batch negatives,
while other hyperparameters remain consistent with the English E5 models~\citep{wang2022text}.

\begin{table}[ht]
\centering
\scalebox{0.95}{\begin{tabular}{lc}
\hline
          & \# Sampled \\ \hline
MS-MARCO Passage &  500k       \\
MS-MARCO Document & 70k \\
NQ, TriviaQA, SQuAD &  220k     \\
NLI  &   275k   \\
ELI5   &  100k      \\
NLLB & 100k \\
DuReader Retrieval  &  86k      \\
Fever &  70k  \\
HotpotQA   &  70k  \\
Quora Duplicate Questions   & 15k   \\
Mr. TyDi   &   50k    \\
MIRACL   &   40k   \\ \hline
Total          &  $\sim$1.6M       \\ \hline
\end{tabular}}
\caption{Data mixture for supervised fine-tuning.}
\label{tab:sup_finetune_data}
\end{table}

\noindent
\textbf{Supervised Fine-tuning }
In the second stage,
we fine-tune the models from the previous stage on a combination of high-quality labeled datasets.
In addition to in-batch negatives,
we also incorporate mined hard negatives and knowledge distillation from a cross-encoder model
to further enhance the embedding quality.
For the \emph{mE5-\{small / base / large\}} models released in mid-2023,
we employ the data mixture shown in Table~\ref{tab:sup_finetune_data}.

For the \emph{mE5-large-instruct} model,
we adopt the data mixture from ~\citet{wang2023improving},
which includes additional $500k$ synthetic data generated by GPT-3.5/4~\citep{OpenAI2023GPT4TR}.
This new mixture encompasses $150k$ unique instructions and covers $93$ languages.
We re-use the instruction templates from ~\citet{wang2023improving}
for both the training and evaluation of this instruction-tuned model.

\section{Experimental Results}

\begin{table}[ht]
\centering
\scalebox{0.95}{\begin{tabular}{lc}
\hline
 & MTEB (56 datasets) \\ \hline
LaBSE & 45.2 \\
Cohere$_\text{multilingual-v3}$ & 64.0 \\
BGE$_\text{large-en-v1.5}$ & 64.2  \\ \hline
mE5$_\text{small}$ &  57.9 \\
mE5$_\text{base}$ & 59.5  \\
mE5$_\text{large}$ & 61.5  \\
mE5$_\text{large-instruct}$ & \textbf{64.4} \\ \hline
\end{tabular}}
\caption{Results on the English portion of the MTEB benchmark.
LaBSE~\citep{feng2022language} is exclusively trained on translation pairs.
Limited information is available regarding the training data and model size are available for Cohere$_\text{multilingual-v3}$(\url{https://txt.cohere.com/introducing-embed-v3/}).
BGE$_\text{large-en-v1.5}$~\citep{xiao2023c} is an English-only model.
Full results are in Appendix Table ~\ref{tab:app_full_results}.}
\label{tab:english_benchmark}
\end{table}

\noindent
\textbf{English Text Embedding Benchmark }
Multilingual embedding models should be able to perform well on English tasks as well.
In Table~\ref{tab:english_benchmark},
we compare our models with other multilingual and English-only models on the MTEB benchmark~\citep{muennighoff2023mteb}.
Our best mE5 model surpasses the previous state-of-the-art multilingual model Cohere$_\text{multilingual-v3}$,
by $0.4$ points and outperforms a strong English-only model,
BGE$_\text{large-en-v1.5}$, by $0.2$ points.
While smaller models demonstrate inferior performance,
their faster inference and reduced storage costs render them advantageous for numerous applications.
\newline

\begin{table}[ht]
\centering
\scalebox{0.95}{\begin{tabular}{lcc}
\hline
 & nDCG@10 & R@100 \\ \hline
BM25 & 39.3 & 78.7  \\
mDPR & 41.5 & 78.8 \\ \hline
mE5$_\text{small}$ &  60.8  & 92.4 \\
mE5$_\text{base}$ &  62.3 &  93.1 \\
mE5$_\text{large}$ & \textbf{66.5} &  94.3 \\
mE5$_\text{large-instruct}$ & 65.7 & \textbf{94.6} \\ \hline
\end{tabular}}
\caption{Multilingual retrieval on the development set of the MIRACL benchmark.
Numbers are averaged over $16$ languages.}
\label{tab:miracl}
\end{table}

\noindent
\textbf{Multilingual Retrieval }
We evaluate the multilingual retrieval capability of our models using the MIRACL benchmark~\citep{Zhang2023MIRACLAM}.
As shown in Table~\ref{tab:miracl},
mE5 models significantly outperform mDPR,
which has been fine-tuned on the MIRACL training set,
in both nDCG@10 and recall metrics.
Detailed results on individual languages are provided in Appendix Table~\ref{tab:app_full_miracl_ndcg}.

\begin{table}[ht]
\centering
\scalebox{0.95}{\begin{tabular}{lcc}
\hline
 & \begin{tabular}[c]{@{}c@{}}BUCC 2018\\ 4 langs\end{tabular} & \begin{tabular}[c]{@{}c@{}}Tatoeba\\ 112 langs\end{tabular} \\ \hline
mContriever$_\text{msmarco}$ & 93.7 & 37.7 \\
LaBSE & 98.8 & 81.1 \\ \hline
mE5$_\text{small}$ & 93.2  & 64.2 \\
mE5$_\text{base}$ & 98.1 &  68.1 \\
mE5$_\text{large}$ & 98.6 &  75.7 \\
mE5$_\text{large-instruct}$ & \textbf{99.0} & \textbf{83.8} \\ \hline
\end{tabular}}
\caption{Bitext mining results.
mContriever~\citep{Izacard2021TowardsUD} numbers are run by ourselves based on the released checkpoint.}
\label{tab:bitext}
\end{table}

\noindent
\textbf{Bitext Mining }
is a cross-lingual similarity search task that requires the matching of two sentences with little lexical overlap.
As demonstrated in Table ~\ref{tab:bitext},
mE5 models exhibit competitive performance across a broad range of languages,
both high-resource and low-resource.
Notably,
the mE5$_\text{large-instruct}$ model surpasses the performance of LaBSE,
a model specifically designed for bitext mining,
due to the expanded language coverage afforded by the synthetic data~\citep{wang2023improving}.

\section{Conclusion}
In this brief technical report,
we introduce multilingual E5 text embedding models that are trained with a multi-stage pipeline.
By making the model weights publicly available,
practitioners can leverage these models for information retrieval, semantic similarity, and clustering tasks
across a diverse range of languages.

\bibliography{anthology,custom}
\bibliographystyle{acl_natbib}

\appendix

\begin{table*}[ht]
\centering
\scalebox{0.9}{\begin{tabular}{lcccccccc}
\hline
& \multicolumn{4}{c}{nDCG@10} & \multicolumn{4}{c}{R@100} \\ \hline
 & mE5$_\text{small}$ & mE5$_\text{base}$ & mE5$_\text{large}$ & E5$_\text{large-instruct}$ & mE5$_\text{small}$ & mE5$_\text{base}$ & mE5$_\text{large}$ & E5$_\text{large-instruct}$ \\ \hline
ar & 71.4 & 71.6 & 76.0 & 76.8 & 96.2 & 95.9 & 97.3 & 97.5 \\
bn & 68.2 & 70.2 & 75.9 & 73.9 & 97.4 & 96.6 & 98.2 & 98.2 \\
en & 48.0 & 51.2 & 52.9 & 51.5 & 85.3 & 86.4 & 87.6 & 88.2 \\
es & 51.2 & 51.5 & 52.9 & 53.7 & 87.6 & 88.6 & 89.1 & 89.3  \\
fa & 53.3 & 57.4 & 59.0 & 59.4 & 90.4 & 91.2 & 92.9 & 92.9  \\
fi & 73.3 & 74.4 & 77.8 & 77.3 & 96.3 & 96.9 & 98.1 & 97.9  \\
fr & 47.6 & 49.7 & 54.5 & 53.7 & 89.5 & 90.0 & 90.6 & 91.7  \\
hi & 55.2 & 58.4 & 62.0 & 60.3 & 91.0 & 92.6 & 93.9 & 94.1  \\
id & 50.7 & 51.1 & 52.9 & 52.1 & 86.2 & 87.4 & 87.9 &  88.4 \\
ja & 63.6 & 64.7 & 70.6 & 69.0 & 95.2 & 96.0 & 97.1 & 96.9  \\
ko & 61.2 & 62.2 & 66.5 & 65.3 & 92.0 & 91.6 & 93.4 & 93.0  \\
ru & 59.1 & 61.5 & 67.4 & 67.9 & 92.2 & 92.7 & 95.5 & 95.4  \\
sw & 68.4 & 71.1 & 74.9 & 72.5 & 94.7 & 95.6 & 96.7 &  97.2 \\
te & 81.3 & 75.2 & 84.6 & 83.4 & 97.6 & 98.0 & 99.2  &  99.0 \\
th & 75.0 & 75.2 & 80.2 & 78.6 & 98.2 & 98.0 & 98.9 &  98.7 \\
zh & 45.9 & 51.5 & 56.0 & 56.2 & 87.9 & 92.1 & 93.3 & 94.9  \\ \hline
Avg & 60.8 & 62.3 & 66.5 & 65.7 & 92.4 & 93.1 & 94.3 & 94.6  \\ \hline
\end{tabular}}
\caption{nDCG@10 and R@100 on the development set of the MIRACL dataset.}
\label{tab:app_full_miracl_ndcg}
\end{table*}

\section{Implementation Details}

\noindent
\textbf{Contrastive Pre-training Text Pairs }
In Table ~\ref{tab:pretrain_data},
to construct text pairs,
we utilize (section title, section passage) for Wikipedia,
(title, page content) for mC4~\citep{xue2021mt5},
(title, news content) for multilingual CCNews~\footnote{\url{https://commoncrawl.org/blog/news-dataset-available}},
translation pairs for NLLB~\citep{costa2022no},
(comment, response) for Reddit~\footnote{\url{https://www.reddit.com/}},
(title, abstract) and citation pairs for S2ORC~\citep{lo2020s2orc},
(question, answer) for Stackexchange~\footnote{\url{https://stackexchange.com/}},
(input prompt, response) for xP3~\citep{muennighoff2022crosslingual}.
For the miscellaneous SBERT data~\footnote{\url{https://huggingface.co/datasets/sentence-transformers/embedding-training-data}},
we include the following datasets:
SimpleWiki, WikiAnswers, AGNews, AltLex, AmazonQA, AmazonReview,
CNN/DailyMail, CodeSearchNet, Flickr30k, GooAQ, NPR, SearchQA,
SentenceCompression, Specter, WikiHow, XSum, and YahooAnswers.
\newline

\noindent
\textbf{Data Mixture for Supervised Fine-tuning }
It includes ELI5~\citep{fan2019eli5}(sample at $20\%$), HotpotQA~\citep{yang2018hotpotqa},
FEVER~\citep{Thorne2018FEVERAL}, MIRACL~\citep{Zhang2023MIRACLAM},
MSMARCO passage ranking and document ranking (sample at $20\%$)~\citep{Campos2016MSMA},
NQ~\citep{Karpukhin2020DensePR}, NLLB (sample at $100k$)~\citep{costa2022no},
NLI~\citep{Gao2021SimCSESC}, SQuAD~\citep{Karpukhin2020DensePR}, TriviaQA~\citep{Karpukhin2020DensePR},
Quora Duplicate Questions~\citep{quora-question-pairs}(sample at $10\%$), MrTyDi~\citep{zhang2021mr},
and DuReader~\citep{qiu2022dureader} datasets.

For the \emph{mE5-large-instruct} model,
we employ the new data mixture from ~\citet{wang2023improving}.
The main difference is the inclusion of synthetic data from GPT-4.
\newline

\noindent
\textbf{Training Hyperparameters }
The mE5$_\text{small}$, mE5$_\text{base}$ and mE5$_\text{large}$
are initialized from the multilingual MiniLM ~\citep{wang2021minilmv2},
\emph{xlm-roberta-base}~\citep{conneau2020unsupervised},
and \emph{xlm-roberta-large} respectively.
For contrastive pre-training,
the learning rate is set to \{$3, 2, 1$\}$\times10^{-4}$ for the \{small, base, large\} models.
For fine-tuning,
we use batch size $512$ and
learning rate \{$3, 2, 1$\}$\times10^{-5}$ for the \{small, base, large\} models.
All models are fine-tuned for $2$ epochs.
The \emph{mE5-large-instruct} model adopts the same hyperparameters as the mE5$_\text{large}$ large,
but is fine-tuned on the new data mixture by ~\citet{wang2023improving}.

\begin{table*}[ht]
\centering
\scalebox{0.88}{\begin{tabular}{lcccc}
\hline
Dataset & mE5$_\text{small}$ & mE5$_\text{base}$ & mE5$_\text{large}$ & mE5$_\text{large-instruct}$ \\ \hline
BIOSSES & 82.3 & 85.1 & 82.5 & 87.0 \\
SICK-R & 77.5 & 78.5 & 80.2 & 81.7 \\
STS12 & 76.6 & 76.7 & 80.0 & 82.6 \\
STS13 & 77.0 & 78.0 & 81.5 & 87.2 \\
STS14 & 75.5 & 76.6 & 77.7 & 85.0 \\
STS15 & 87.1 & 88.2 & 89.3 & 91.0 \\
STS16 & 83.6 & 84.3 & 85.8 & 87.3 \\
STS17 & 86.4 & 87.8 & 88.1 & 90.0 \\
STS22 & 60.9 & 61.8 & 63.1 & 67.6 \\
STSBenchmark & 84.0 & 85.6 & 87.3 & 88.4 \\
SummEval & 30.0 & 30.1 & 29.7 & 30.4 \\
SprintDuplicateQuestions & 92.2 & 93.0 & 93.1 & 91.2 \\
TwitterSemEval2015 & 70.8 & 72.2 & 75.3 & 80.3 \\
TwitterURLCorpus & 84.8 & 85.5 & 85.8 & 87.1 \\
AmazonCounterfactualClassification & 73.8 & 79.0 & 79.1 & 76.2 \\
AmazonPolarityClassification & 88.7 & 90.6 & 93.5 & 96.3 \\
AmazonReviewsClassification & 44.7 & 44.5 & 47.6 & 56.7 \\
Banking77Classification & 79.4 & 82.7 & 84.7 & 85.7 \\
EmotionClassification & 42.5 & 45.2 & 46.5 & 51.5 \\
ImdbClassification & 80.8 & 85.5 & 90.2 & 94.6 \\
MassiveIntentClassification & 70.3 & 72.1 & 73.8 & 77.1 \\
MassiveScenarioClassification & 74.5 & 77.1 & 77.5 & 80.5 \\
MTOPDomainClassification & 91.1 & 93.1 & 93.7 & 93.9 \\
MTOPIntentClassification & 71.1 & 75.3 & 77.9 & 82.5 \\
ToxicConversationsClassification & 69.4 & 69.8 & 71.3 & 71.1 \\
TweetSentimentExtractionClassification & 62.6 & 61.3 & 62.0 & 64.6 \\
AskUbuntuDupQuestions & 57.9 & 58.2 & 60.3 & 63.9 \\
MindSmallReranking & 30.3 & 31.0 & 31.4 & 33.1 \\
SciDocsRR & 78.1 & 80.7 & 82.0 & 85.9 \\
StackOverflowDupQuestions & 49.2 & 49.4 & 49.7 & 51.5 \\
ArxivClusteringP2P & 39.2 & 40.3 & 44.3 & 46.4 \\
ArxivClusteringS2S & 30.8 & 35.4 & 38.4 & 40.5 \\
BiorxivClusteringP2P & 35.8 & 35.0 & 35.3 & 40.9 \\
BiorxivClusteringS2S & 27.1 & 29.5 & 33.5 & 36.3 \\
MedrxivClusteringP2P & 30.9 & 28.9 & 31.5 & 36.9 \\
MedrxivClusteringS2S & 27.3 & 28.4 & 29.7 & 35.5 \\
RedditClustering & 39.1 & 42.4 & 46.5 & 56.6 \\
RedditClusteringP2P & 59.0 & 55.2 & 63.2 & 64.3 \\
StackExchangeClustering & 53.5 & 55.3 & 57.5 & 66.8 \\
StackExchangeClusteringP2P & 32.1 & 30.5 & 32.7 & 42.5 \\
TwentyNewsgroupsClustering & 33.2 & 36.0 & 38.9 & 51.3 \\
ArguAna & 39.1 & 44.2 & 54.4 & 58.4 \\
ClimateFEVER & 22.6 & 23.9 & 25.7 & 29.9 \\
CQADupstackAndroidRetrieval & 36.1 & 38.5 & 39.7 & 42.7 \\
DBPedia & 37.8 & 40.4 & 41.3 & 38.4 \\
FEVER & 75.3 & 79.4 & 82.8 & 78.0 \\
FiQA2018 & 33.3 & 38.2 & 43.8 & 47.7 \\
HotpotQA & 65.1 & 68.6 & 71.2 & 69.3 \\
MSMARCO & 41.0 & 42.3 & 43.7 & 40.4 \\
NFCorpus & 31.0 & 32.5 & 34.0 & 35.5 \\
NQ & 56.3 & 60.0 & 64.1 & 57.8 \\
QuoraRetrieval & 86.9 & 87.7 & 88.2 & 89.2 \\
SCIDOCS & 13.9 & 17.2 & 17.5 & 18.7 \\
SciFact & 67.7 & 69.3 & 70.4 & 71.8 \\
Touche2020 & 21.2 & 21.4 & 23.4 & 27.2 \\
TRECCOVID & 72.6 & 69.8 & 71.3 & 82.0 \\ \hline
Average & 57.9 & 59.4 & 61.5 & \textbf{64.4} \\ \hline
\end{tabular}}
\caption{Results for each dataset in the MTEB benchmark.
The evaluation metrics are available in the original paper~\citep{muennighoff2023mteb}.}
\label{tab:app_full_results}
\end{table*}

\end{document}